morganislame:

Solve 5x+2(x+2)<-10 (meant to be less than or equal to)

Ultrilliam:

\(5x+2\times(x+2)\le-10\)?

morganislame:

Yes, im sorry for not writing it correctly.

Vocaloid:

first distribute the 2 in 2(x+2)

Vocaloid:

A(B+C) = A*B + A*C 2(x+2) = ?

morganislame:

2x + 4?

Vocaloid:

awesome! so that gives us 5x+2x + 4<= -10 now we add the 5x + 2x to get 7x + 4 <= -10

Vocaloid:

then we subtract 4 from both sides 7x <= -14 then divide both sides by 7 to get your final answer

morganislame:

So that means -2 is the answer, right?

Vocaloid:

good, x <= -2 is your answer (supposed to be a less than or equal sign)

morganislame:

Okay, that makes sense now. Thank you sm! :)

Vocaloid:

imtoolazytouselatexhushultri

Ultrilliam:

vocaloid \gt = greater then \lt = less then HA YOU KNEW WHAT I WAS TYPING X'D

Ultrilliam:

\ge = greater then or equal to \lt = less then or equal to